OpenStudy (anonymous):

The graph of which quadratic equation has a vertex of (–2, –1)?

OpenStudy (anonymous):

A.y = (x – 2)2 – 1 B.y = (x + 1)2 – 2 C.y = (x + 2)2 – 1 D.y = (x – 1)2 – 2

OpenStudy (anonymous):

\[ax^2+bx+c\] \[(\frac{-b}{2a},f(\frac{-b}{2a}))\] Looking at your options, choice C looks most likely for -b/2a to = -2 Expand: \[x^2+4x+4-1=x^2+4x+3\] -b/(2a)=-4/2(1)=-2 \[f(-2)=(-2)^2+4(-2)+3=4-4+3=-1\] Therefore; option C is the answer.